4(3y-1)+13=5y+2
Answer is..Y = ?

Answer by jb007mi6(53)

this is simple
first i would distribute the 4 so you get
12y-4+13=5y=2
then i would simplify the left side
12y+9=5y-2
then I would subtract 9 to both sides
12y=5y-11
then i would subtract -5 to both sides
7y=-11
divide by 7
y=-11/7